# InstantSplat: Sparse-view Gaussian Splatting in Seconds

### v5

@@ 3.2 Co-visible Global Geometry Initialization
With the off-the-shell stereo model, MASt3R, we regress a set of stereo point maps along with calculated camera intrinsics and extrinsics following Sec. 3.1. To align these pair wise predictions into a globally aligned results, we follow MASt3R to build a complete connectivity graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ of all the N input views, where the vertices $\mathcal{V}$ and each edge $e=(I_{i},I_{j})\in\mathcal{E}$ indicate a shared visual content between images $I_{i}$ and $I_{j}$ ($i,j\in\left\{1,2,\cdots,N\right\},i\neq j$). To convert the initially predicted point map $\{(\mathbf{P}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1}^{N}$ to be a globally aligned one $\{(\mathbf{\tilde{P}}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1}^{N}$, we follow the MASt3R to update these point maps, transformation matrix, and a scale factor into globally corrdinate system. Please refer to the supplementary material for detailed description. However, the pairwise predicted focal lengths remain inconsistent post-optimization. Assuming a single-camera setup akin to COLMAP for capturing a scene, we stabilize the estimated focal lengths by averaging them across all training views: $\bar{f}=\frac{1}{N}\sum_{i=1}^{N}f_{i}^{*}$ which is simple but significantly improve optimized visual quality.
@@
$s_{i}=\frac{1}{|\mathbf{O}_{i}|}\sum_{o\in\mathbf{O}_{i}}o$ (5)
where $s_{i}$ denotes the average confidence score of view $i$.
Co-visible redundancy elimination starts with the view ($j$) having the lowest average confidence score, where we combine point maps from views with higher average confidence score than view $j$, $\{(\mathbf{\tilde{P}}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1,i\neq j,s_{i}>s_{j}}^{N}$, and project them onto view $j$. Depth ambiguity is addressed through a depth verification process, pruning points only if the difference between the projected and original depth values is higher than a predefined threshold: $\Delta d=|d_{\text{proj}}-d_{\text{orig}}|$, where $d_{\text{orig}}$ are obtained from MASt3R. This procedure is iterated in ascending order across all images, ultimately generating binary masks $\{(\mathcal{M}_{i}\in\mathbb{R}^{H\times W\times 1})\}_{i=1}^{N}$ to filter out redundant points. More formally, the cross-view co-visibility check for removing redundancy in view $j$ can be written as:
$\mathbf{D}_{\text{proj},j}$ =$=\bigcap_{\{i\mid s_{i}>s_{j}\}}\text{Proj}_{j}(\mathbf{\tilde{P}}_{i})$ (6)
@@
$\mathbf{\tilde{P}}_{j}$ $=(1-\mathcal{M}_{j})\cdot\mathbf{\tilde{P}}_{j}$ (8)
@@ 3.3 Rapid Optimization
Joint Optimization for Alignment
Given multiple views and a 3D representation with initialized poses, represented by a set of Gaussians $\mathbf{G}$, and camera poses $\mathbf{T}$, we employ the pre-built rasterizer from Gaussian Splatting as a differentiable operator to render the images corresponding to the target poses $\mathbf{T}$. Gradient descent with photometric error minimization is used to reduce the discrepancy between the optimizable 3D representation and 2D observations, iteratively refining $\mathbf{G}$, and $\mathbf{T}$ toward representative orientations:
$\footnotesize\mathbf{G}^{*},\mathbf{T}^{*}=\underset{\mathbf{G},\mathbf{T}}{\arg\min}\sum_{v\in N}\sum_{i=1}^{HW}$ $\left\|\tilde{\mathbf{C}}_{v}^{i}-\mathbf{C}_{v}^{i}(\mathbf{G},\mathbf{T})\right\|$
Aligning Camera Poses on Test Views
Unlike benchmark datasets where camera pose estimation and calibration are pre-computed with access to all training and testing views, a COLMAP-free setting requires additional steps to regress the poses for the test views. Similar to INeRF [41], we freeze the 3D model learned by training images, while optimizing the camera poses for the test views. This optimization process focuses on minimizing the photometric discrepancies between the synthesized images and the actual images on test views, aiming to achieve alignment between the 3D model and test images for evaluating visual quality and pose estimation metrics.
@@
Table 2: Quantitative Evaluations on Tanks and Temples Datasets. Our method achieves significantly clearer details (lower LPIPS) compared to other COLMAP-free methods, even with an optimization time of only $\sim$7.5 seconds (Ours-S), and avoids artifacts caused by noisy pose estimation, as quantified by Absolute Trajectory Error (ATE). Extending training time further improves rendering quality. InstantSplat also demonstrates superior camera pose estimation accuracy in ATE with ground truth scale. The optimization iterations is set 200 for Ours-S, and 1,000 for Ours-XL.
@@
Figure 4: Visual Comparisons between InstantSplat and various COLMAP-free methods. InstantSplat achieves faithful 3D reconstruction and renders novel views with as few as three training images on the Tanks and Temples datasets, and the MVImgNet datasets.
@@ 4.1 Experimental Setup
Table 3: Quantitative Evaluations on MVImgNet Datasets. Our method renders significantly clearer details (by LPIPS) compared to other methods, devoid of artifacts typically associated with noisy pose estimation (e.g., CF-3DGS [12], NeRFmm [41] and NoPe-NeRF [5]). COLMAP on full views as the ground-truth reference. Our method produces more accurate camera pose estimation in the Absolute Trajectory Error (ATE) than previous COLMAP-free methods, quantified using the ground truth scale.
@@
NoPe-NeRF [5] 37min42s 53min 95min 0.4326 0.4329 0.4686 0.6168 0.6614 0.6257 0.2780 0.1740 0.1493
CF-3DGS [12] 3min47s 7min29s 10min36s 0.3414 0.3544 0.3655 0.4520 0.4326 0.4492 0.1593 0.1981 0.1243
NeRF-mm [41] 8min35s 16min30s 33min14s 0.3752 0.3685 0.3718 0.6421 0.6252 0.6020 0.2721 0.2376 0.1529
@@
Following methodologies from previous research [12, 5], we evaluated our InstantSplat on eight scenes from the Tanks and Temples dataset [21] with view counts ranging from 3 to 12. Additionally, we tested on seven outdoor scenes from the MVImgNet dataset [50], covering diverse scene types including Car, Suv, Bicycle, Chair, Ladder, Bench, and Table. Our experiments also extend to in-the-wild data, incorporating frames from the Sora video (link), and navigation stereo camera data from the Perseverance Rover (NASA), along with three randomly sampled training views from the DL3DV-10K dataset [25].
@@
Our comparisons on COLMAP-free methods include Nope-NeRF [5] and CF-3DGS [12], both supported by monocular depth maps and ground-truth camera intrinsics. We also consider NeRFmm [41], which involves joint optimization of NeRF and all camera parameters. Additionally, we compare our method with 3D-GS [20] and FSGS [55], which utilize COLMAP for pre-computing the camera parameters.
@@
Table 4: Ablation Study for validating design choices. We opt for the XL setting, where 1,000 iterations are adopted to ensure reconstruction quality.
@@ 4.3 Ablations and Analysis
We conducted ablation studies to validate our design choices, transitioning from the use of non-differentiable Structure-from-Motion (SfM) complemented by Gaussian Splatting with adaptive density control, to adopting Multi-View Stereo (MVS) with Co-visible Global Geometry Initialization, Confidence-aware Optimizer and Joint Optimization for efficient and robust sparse-view 3D modeling. Experiments were conducted using 12 training views on all scenes in the Tanks and Temples dataset, with COLMAP poses derived from full dense views serving as ground truth, unless otherwise specified.
@@
  • Question 4: Is Joint Optimization essential for accurate rendering?
@@
For Q1, the evaluation of camera pose accuracy from MASt3R reveals considerable differences in rendering quality and pose accuracy, as shown in the first row of Tab. 4. This baseline is implemented by directly stitching together MASt3R’s point cloud and camera parameters for the initialization of 3D-GS to optimize the Gaussian attributes.
@@
For Q3, co-visible geometry initialization significantly enhances FPS by minimizing redundancy while preserving robust rendering accuracy, as elaborated in the third row.
For Q4, Joint Optimization, through self-supervised optimization, substantially improves visual and pose metrics by aligning the 3D and 2D representations.
@@
Table 6: InstantSplat with 2D-GS. InstantSplat demonstrates strong compatibility with 2D-GS, notably improving sparse-view rendering quality and depth accuracy while also accelerating scene reconstruction. We opt for 1,500 optimization iterations to balance reconstruction quality and efficientcy.
@@ 5 Conclusion
We present InstantSplat, the first system that synergizes the geometric priors from pre-trained deep models, while performing joint optimization for accurate 3D scene reconstruction and camera pose estimation. InstantSplat is efficient and effective under sparse-view reconstruction, capable of building a 3D model in just a few seconds, thereby demonstrating superior practicality in real-world scenarios. It aligns the 3D representation and 2D images with a differentiable operator and avoids the time-consuming and fragile Structure-from-Motion and Adaptive Density Control. InstantSplat also introduces co-visibility-based initialization as well as confidence-aware optimization to enhance robustness. Compared with previous best-performing COLMAP-free methods [12, 5], InstantSplat dramatically reduces the required number of views from hundreds to just a few and demonstrates robustness across in-the-wild datasets, as well as adaptability to many point-based representations.
@@
However, InstantSplat faces limitations when applied to scenes with more than a few hundred images due to the pixel-aligned initialization. Addressing this issue through progressive alignment with Gaussian pruning techniques will be the focus of future research.

### v6

@@ 3.3 Co-visible Global Geometry Initialization
With the off-the-shell stereo model, MASt3R, we regress a set of stereo point maps along with calculated camera intrinsics and extrinsics following Sec. 3.1. To align these pair wise predictions into a globally aligned results, we follow MASt3R to build a complete connectivity graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ of all the N input views, where the vertices $\mathcal{V}$ and each edge $e=(n,m)\in\mathcal{E}$ indicate a shared visual content between images $I_{n}$ and $I_{m}$. To convert the initially predicted point map $\{(\mathbf{P}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1}^{N}$ to be a globally aligned one $\{(\mathbf{\tilde{P}}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1}^{N}$, we follow the MASt3R to update these point maps, transformation matrix, and a scale factor into globally corrdinate system. Please refer to the supplementary material for detailed description. However, the pairwise predicted focal lengths remain inconsistent post-optimization. Assuming a single-camera setup akin to COLMAP for capturing a scene, we stabilize the estimated focal lengths by averaging them across all training views: $\bar{f}=\frac{1}{N}\sum_{i=1}^{N}f_{i}^{*}$ which is simple but significantly improve optimized visual quality.
@@
$v_{i}=\frac{1}{|\mathbf{O}_{i}|}\sum_{o\in\mathbf{O}_{i}}o$ (5)
where $v_{i}$ denotes the average confidence score of view $i$.
Cross-view checking starts with the view ($j$) having the lowest average confidence, where we combine point maps from views with higher average confidence score than view $j$, $\{(\mathbf{\tilde{P}}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1,i\neq j,v_{i}>v_{j}}^{N}$, and project them onto view $j$. Depth ambiguity is addressed through a depth verification process, pruning points only if the difference between the projected and original depth values is higher than a predefined threshold: $\Delta d=|d_{\text{proj}}-d_{\text{orig}}|$. This procedure is iterated in ascending order across all images, ultimately generating a binary mask $\{(\mathbf{\tilde{M}}_{i}\in\mathbb{R}^{H\times W\times 1})\}_{i=1}^{N}$ to filter out redundant points. More formally, the cross-view co-visibility check for removing redundancy in view $j$ can be written as:
$\mathbf{D}_{\text{proj},j}$ =$=\bigcap_{\{i\mid v_{i}>v_{j}\}}\text{Proj}_{j}(\mathbf{\tilde{P}}_{i})$ (6)
@@
$\mathbf{P}_{j}$ $=(1-\mathcal{M}_{j})\cdot\mathbf{\tilde{P}}_{j}$ (8)
@@ 3.4 Optimization
Training Objective
To optimize our neural 3D representation efficiently, we employ purely visual signals, using photometric loss to minimize discrepancies between the rasterized images from our 3D model, denoted as $\tilde{\mathbf{C}}(\mathbf{G},\mathbf{T})$, and the observed images $\mathbf{C}$. Overall, the process maps observed unposed RGB images, to a 3D representation $\mathbf{G^{*}}$ with accurate pose estimations $\mathbf{T^{*}}$, utilizing end-to-end iterative optimization and a self-supervised training objective. InstantSplat aligns scene representations and camera poses toward visual observation, improving rendering accuracy and optimization efficiency without additional heuristic regularization.
@@
mechanism efficiently adapts parameters through the use of photometric loss, applicable to both camera and Gaussian parameters.
Table 2: Quantitative Evaluations on Tanks and Temples Datasets. Our method achieves significantly clearer details (lower LPIPS) compared to other pose-free methods, even with an optimization time of only $\sim$7.5 seconds (Ours-S), and avoids artifacts caused by noisy pose estimation, as quantified by Absolute Trajectory Error (ATE). Extending training time further improves rendering quality. InstantSplat also demonstrates superior camera pose estimation accuracy in ATE with ground truth scale. In 3D-GS, ADC refers to Adaptive Density Control [21]. We further leverage cupy to accelerate MASt3R inference.
@@
Figure 4: Visual Comparisons between InstantSplat and various pose-free baseline methods. InstantSplat achieves faithful 3D reconstruction and renders novel views with as few as three training images on the Tanks and Temples datasets, and the MVImgNet datasets.
@@ 4.1 Experimental Setup
Table 3: Quantitative Evaluations on MVImgNet Datasets. Our method renders significantly clearer details (by LPIPS) compared to other pose-free methods, devoid of artifacts typically associated with noisy pose estimation (e.g., CF-3DGS [13], NeRFmm [42]) and NoPe-NeRF [5]. COLMAP on full views as the ground-truth reference. Our method produces more accurate camera pose estimation in the Absolute Trajectory Error (ATE) than previous COLMAP-free methods, quantified using the ground truth scale. ADC is denoted as Adaptive Density Control [21]. Cupy library is used to accelerate the initialization.
@@
NoPe-NeRF [5] 37min42s 53min 95min 0.4326 0.4329 0.4686 0.6674 0.6614 0.6257 0.2780 0.1740 0.1493
CF-3DGS [13] 3min47s 7min29s 10min36s 0.3414 0.3544 0.3655 0.5523 0.4326 0.4492 0.1593 0.1981 0.1243
NeRF-mm [42] 8min35s 16min30s 33min14s 0.3752 0.3685 0.3718 0.7001 0.6252 0.6020 0.2721 0.2376 0.1529
@@
Following methodologies from previous pose-free research [13, 5], we evaluated our InstantSplat on eight scenes from the Tanks and Temples dataset [22] with view counts ranging from 3 to 12. Additionally, we tested on seven outdoor scenes from the MVImgNet dataset [52], covering diverse scene types including Car, Suv, Bicycle, Chair, Ladder, Bench, and Table. Our experiments also extend to in-the-wild data, incorporating frames from the Sora video (link), and navigation stereo camera data from the Perseverance Rover (NASA), along with three randomly sampled training views from the DL3DV-10K dataset [26].
@@
Our comparisons on pose-free methods include Nope-NeRF [5] and CF-3DGS [13], both supported by monocular depth maps and ground-truth camera intrinsics. We also consider NeRFmm [42], which involves joint optimization of NeRF and all camera parameters. Additionally, we compare our method with 3D-GS [21] and FSGS [57], which utilize COLMAP for pre-computing the camera parameters.
@@
Table 4: Ablation Study for validating design choices. We opt for the XL setting, where 1,000 iterations are adopted to ensure reconstruction quality. Results are averaged on all scenes of Tank and Temple datasets while Cupy library is adopted.
@@ 4.3 Ablations and Analysis
We conducted ablation studies to validate our design choices, transitioning from the use of non-differentiable Structure-from-Motion (SfM) complemented by Gaussian Splatting with adaptive density control, to adopting Multi-View Stereo (MVS) with Co-visible Global Geometry Initialization, Confidence-aware Optimizer and Gaussian Bundle Adjustment for efficient and robust sparse-view 3D modeling. Experiments were conducted using 12 training views on all scenes in the Tanks and Temples dataset, with COLMAP poses derived from full dense views serving as ground truth, unless otherwise specified.
@@
  • Question 4: Is Gaussian Bundle Adjustment essential for accurate rendering?
@@
For Q1, the evaluation of camera pose accuracy from MASt3R reveals significant discrepancies in rendering quality and pose accuracy, as shown in Tab. 4’s first and last rows.
@@
For Q3, co-visible geometry initialization markedly improves FPS by reducing redundancy. Confidence-based view ranking further enhances rendering accuracy, as detailed in the third and fourth rows.
For Q4, Gaussian Bundle Adjustment (GauBA), through self-supervised optimization, substantially improves visual and pose metrics by aligning the 3D and 2D representations.
@@
Table 6: InstantSplat with 2D-GS. InstantSplat demonstrates strong compatibility with 2D-GS, notably improving sparse-view rendering quality and depth accuracy while also accelerating training by eliminating the need for an additional SfM step. Cupy library is used to accelerate the initialization.
@@ 5 Conclusion
InstantSplat is a system tailored to reconstruct large-scale scenes from sparse-view, unposed images. Operating within a self-supervised framework, InstantSplat optimizes camera poses and 3D representations using photometric signals and avoids the time-consuming Adaptive Density Control. The system leverages multi-view stereo priors for surface point initialization and employs co-visible cross-view checking to minimize model redundancy. Compared with previous best-performing pose-free methods [13, 5], InstantSplat dramatically reduces the required number of views from hundreds to just a few and demonstrates robustness across in-the-wild datasets, as well as adaptability to various point-based representations.
@@
However, InstantSplat faces limitations when applied to scenes with more than hundreds of images due to its reliance on MVS for globally aligned point clouds. Addressing this through progressive alignment techniques will be a focus of future research.
A6 Details of Global Geometry Initialization
We provide additional details on the scaling process for pairwise-to-global alignment as discussed in Co-visible Global Geometry Initialization (Section 3.3 of the main draft).
Specifically, the geometric prior model, MASt3R, originally uses image pairs as input, requiring post-processing to align scales when more than two views are captured from the scene. This necessity arises because the predicted point maps are generated at varying scales. Such scale misalignment across independently computed relative poses introduces variance, leading to inaccuracies in camera pose estimation.
To integrate these pairwise pointmaps and camera information into a globally aligned geometry, we follow MASt3R [23] to construct a complete connectivity graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ of all the N input views, where the vertices $\mathcal{V}$ and each edge $e=(n,m)\in\mathcal{E}$ indicate a shared visual content between images $I_{n}$ and $I_{m}$. To convert the initially predicted point map $\{(\bm{P}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1}^{N}$ to be a globally aligned one $\{(\bm{\tilde{P}}_{i}\in\mathbb{R}^{H\times W\times 3})\}_{i=1}^{N}$, we update the point maps, transformation matrix, and a scale factor: For the complete graph, any image pair $e=(n,m)\in\mathcal{E}$, the point maps $\bm{P}_{n,n}$,$\bm{P}_{m,n}$ and confidence maps $\bm{O}_{n,n}$,$\bm{O}_{m,n}$. For clarity, let us define $\bm{P}_{n,e}:=\bm{P}_{n,n}$, and $\bm{P}_{m,e}:=\bm{P}_{m,n}$, $\bm{O}_{n,e}:=\bm{O}_{n,n}$, and $\bm{O}_{m,e}:=\bm{O}_{m,n}$. The optimization for the transformation matrix of edge $\bm{T}_{e}$, scaling factor $\sigma_{e}$ and point map $\tilde{\bm{P}}$ are given by:
$\tilde{\bm{P}}^{*}=\underset{\tilde{\bm{P}},\bm{T},\sigma}{\arg\min}\sum_{e\in\mathcal{E}}\sum_{v\in e}\sum_{i=1}^{HW}\bm{O}_{v,e}^{i}\left\|\tilde{\bm{P}}_{v}^{i}-\sigma_{e}\bm{T}_{e}\bm{P}_{v,e}^{i}\right\|.$ (9)
Here, we slightly abuse notation and use $v\in e$ for $v\in\{n,m\}$ if $e=(n,m)$. The idea is that, for a given pair $e$, the same rigid transformation $\bm{T}_{e}$ should align both pointmaps $\bm{P}_{n,e}$ and $\bm{P}_{m,e}$ with the world-coordinate pointmaps $\tilde{\bm{P}}_{n,e}$ and $\tilde{\bm{P}}_{m,e}$, since $\bm{P}_{n,e}$ and $\bm{P}_{m,e}$ are by definition both expressed in the same coordinate frame. To avoid the trivial optimum where $\sigma_{e}=0,\,\forall e\in\mathcal{E}$, we enforce $\prod_{e}\sigma_{e}=1$. Having aligned the point clouds, we commence the integration process by initializing 3D Gaussians [21], treating each point as a primitive. This post-processing step yields a globally aligned 3D Gaussians within only seconds, where inferring per-view point and confidence maps can be achieved real-time on a modern GPU.
A6.1 Additional Implementation Details.
We implement InstantSplat-2D-GS using the original 2D-GS CUDA kernel. All core components of our system are integrated into 2D-GS, including Focal Averaging, co-visible Global Geometry Initialization, Confidence-aware Optimization, and Gaussian Bundle Adjustment. The optimization iterations are set to 1,500 for Gaussian training and 500 for aligning test view camera poses. Regularization for depth distortion and normal consistency begins at iterations 500 and 700, respectively.
A6.2 Evaluation on 2D-GS
In Table 6, we present the performance of InstantSplat-2D-GS on the tasks of novel view synthesis and multi-view stereo depth estimation using the Tanks and Temples dataset [22]. Following the train/test dataset split scheme described in the main draft, we sample 12 images for training and 12 for evaluation, covering the entire dataset. The test set contains 12 uniformly selected images (excluding first/last frames), with an equal number of training images sampled from remaining frames for sparse-view training. For novel view synthesis, we evaluate performance using PSNR, SSIM, and LPIPS. For depth estimation, we use the Absolute Relative Error (Rel) and Inlier Ratio ($\tau$) with a threshold of 1.03 to assess each scene, following the approach of DUSt3R [40]. We align the scene scale between the predictions and the ground truth, as our system does not depend on camera parameters for prediction. We normalize the predicted depth maps using the median of the predicted depths and similarly normalize the ground truth depths, which follows the procedures established in previous studies [34, 40] to ensure alignment between the predicted and ground truth depth maps. We use COLMAP [33] to perform dense stereo reconstruction on the entire video sequence, generating a high-quality multi-view stereo point cloud. The 3D dense points are then projected onto 2D image planes to create pseudo ground-truth depth maps for evaluation.
A7 More Experimental Results
A7.1 Additional Qualitative Results on the Benchmarks
We provide more visualized results and pose estimations on realistic outdoor scenes from the Tanks and Temples dataset [22] and MVImgNet [52]. As shown in Fig.A9 and Fig.A7, InstantSplat consistently produces fewer artifacts than Gaussian-based methods (e.g., CF-3DGS [13]) and exhibits sharper details than NeRF-based methods (e.g., Nope-NeRF [5] and NeRFmm [42]).
A7.2 Video Demonstration
We provide free-viewpoint rendering on various in-the-wild datasets, including the Sora video, by uniformly sampling three frames from the entire video sequence. Additionally, we use stereo camera data from the Perseverance Rover, available on the NASA website, and randomly sample three training views from the DL3DV-10K dataset [26]. As shown in Fig. A8 and Fig. A9, InstantSplat achieves accurate 3D reconstruction and high-quality novel view synthesis with as few as three training images, highlighting the robustness and effectiveness of our framework. For additional visualizations, please refer to our instantsplat-webpage.zip with opening the “index.html”.
Figure A5: Additional Visual Comparisons between 2D-GS and InstantSplat-2D-GS. We present the rendering results of RGB images, depth maps, and normal maps from the Tanks and Temples dataset. The corresponding SSIM, LPIPS, and training time metrics are displayed on the right. Here, IS is used as an abbreviation for InstantSplat.
Figure A6: Additional Visual Comparisons on the Tanks and Temples Dataset. We present the rendering results from NeRFmm, Nope-NeRF, CF-3DGS, and Ours-XL.
Figure A7: Additional Visual Comparisons on the MVimgNet Dataset. We present the rendering results from NeRFmm, Nope-NeRF, CF-3DGS, and Ours-XL.
Figure A8: Free-view Rendering on the Tanks and Temples Dataset.
Figure A9: Free-view Rendering on in-the-wild data. We present the rendering results of three different in-the-wild data: Mars Rover Navigation, Sora Video and DL3DV-10K dataset